# Scaling Up to Excellence: Practicing Model Scaling for Photo-Realistic Image Restoration In the Wild

### v1

@@ 4.4 Ablation Study
The proposed restoration-guided sampling method is mainly controlled by the hyper-parameter $\tau_{r}$. The larger $\tau_{r}$ is, the fewer corrections are made to the generation at each step. The smaller $\tau_{r}$ is, the more generated content will be forced to be closer to the LQ image. Please refer to Fig. 13 for a qualitative comparison. When $\tau_{r}=0.5$, the image is blurry because its output is limited by the LQ image and cannot generate texture and details. When $\tau_{r}=6$, there is not much guidance during generation. The model generates a lot of texture that is not present in the LQ image, especially in flat area. Fig. 8(a) illustrates the quantitative results of restoration as a function of the variable $\tau_{r}$. As shown in Fig. 8(a), decreasing $\tau_{r}$ from 6 to 4 does not result in a significant decline in visual quality, while fidelity performance improves. As restoration guidance continues to strengthen, although PSNR continues to improve, the images gradually become blurry with loss of details, as depicted in Fig. 13. Therefore, we choose $\tau_{r}=4$ as the default parameter, as it doesn’t significantly compromise image quality while effectively enhancing fidelity.
@@ A.4 Negative Samples Generation
While negative prompts are highly effective in enhancing quality, the lack of negative-quality samples and prompts in the training data results in the fine-tuned SUPIR’s inability to comprehend these prompts effectively. To address this problem, in Sec. 3.2 of the main text, we introduce a method to distill negative concepts from the SDXL model. The process for generating negative samples is illustrated in Fig. 19. Direct sampling of negative samples through a text-to-image approach often results in meaningless images. To address this issue, we also utilize training samples from our dataset as source images. We create negative samples in an image-to-image manner as proposed in [56], with a strength setting of 0.50.50.5.

### v2

@@ 4.4 Ablation Study
The proposed restoration-guided sampling method is mainly controlled by $\tau_{r}$. The larger $\tau_{r}$ is, the fewer corrections are made to the generation at each step. The smaller $\tau_{r}$ is, the more generated content will be forced to be closer to the LQ image. Please refer to Fig. 13 for a qualitative comparison. When $\tau_{r}=0.5$, the image is blurry because its output is limited by the LQ image and cannot generate texture and details. When $\tau_{r}=6$, there is not much guidance during generation. The model generates a lot of texture that is not present in the LQ image, especially in flat area. Fig. 8(a) illustrates the quantitative results of restoration as a function of the variable $\tau_{r}$. As shown in Fig. 8(a), decreasing $\tau_{r}$ from 6 to 4 does not result in a significant decline in visual quality, while fidelity performance improves. As restoration guidance continues to strengthen, although PSNR continues to improve, the images gradually become blurry with loss of details, as depicted in Fig. 13. Therefore, we choose $\tau_{r}=4$ as the default parameter, as it doesn’t compromise image quality while effectively enhancing fidelity.
@@ 5 Conclusion
Acknowledgment.
This work was supported in part by the National Natural Science Foundation of China (62276251, 62272450), the Joint Lab of CAS-HK, the National Key R&D Program of China (No. 2022ZD0160100), and in part by the Youth Innovation Promotion Association of Chinese Academy of Sciences (No. 2020356).
@@ A.4 Negative Samples Generation
While negative prompts are highly effective in enhancing quality, the lack of negative-quality samples and prompts in the training data results in the fine-tuned SUPIR’s inability to comprehend these prompts effectively. To address this problem, in Sec. 3.2 of the main text, we introduce a method to distill negative concepts from the SDXL model. The process for generating negative samples is illustrated in Fig. 19. Direct sampling of negative samples through a text-to-image approach often results in meaningless images. To address this issue, we also utilize training samples from our dataset as source images. We create negative samples in an image-to-image manner as proposed in [56], with a strength setting of $0.5$.